OpenStudy (anonymous):

If f(x) =1/3 x − 2 and g(x) = −2x − 1, find (f • g)(9)

OpenStudy (helder_edwin):

it multiplication no composition i guess. so by definition \[ \large (f\cdot g)(9)=f(9)\cdot g(9)= \]

OpenStudy (anonymous):

is your question the composition of f(g(x)) ???